OpenStudy (anonymous):

Deal with the ambiguous SSA case. For each find all possible solutions. 1. A=50 degrees, a=15, b=12 2. A=50 degreees, a=10, b=15 3. A=47 degrees, a=1.5 b=2

OpenStudy (anonymous):

I just want to know if number two is possible...

OpenStudy (anonymous):

When I use the law of sines, I end up getting sin^-1(B) is undefined. (for number 2)

OpenStudy (luigi0210):

Well if you did it right, plug it into a calculator and get undefined then it's probably not possible.

OpenStudy (anonymous):

can you check my work? sin(B)/15=sin(50)/10 15sin(50)/10=1.14 sin^-1(1.14)=undefined

OpenStudy (luigi0210):

Looks right to me, sine cannot be >1

OpenStudy (anonymous):

Okay thanks @Luigi0210

OpenStudy (luigi0210):

No Problem :)